OpenStudy (anonymous):

Simplify each expression. Use positive exponents. M ^{3}N ^{-6}P^{0} I was never good at math and the book I have is no use at all. I think the n ^–6 turns out to be 1/n^6 .... I think. Any help to understand it would be wonderful.

OpenStudy (asnaseer):

you are correct with n^(-6) = 1/n^6 this site may be helpful to you: http://www.mathsisfun.com/exponent.html let me know if you are still stuck after looking at this site.

OpenStudy (anonymous):

I looked at the site got all the questions it asked right. It's just they are using numbers and not letters. I think that's why I am more confused. I know \[n ^{-6}\] is \[\frac{ 1 }{ n ^{6} } \] and \[p ^{0} \] is 1 I believe. Other then that I am lost.

OpenStudy (asnaseer):

which bit exactly are you lost on?

OpenStudy (asnaseer):

if you use the results you gave above and put them into your original expression then you should end up with the correct answer

OpenStudy (asnaseer):

\[m^3n^{-6}p^0=...\]

OpenStudy (anonymous):

So it's \[m ^{3}\frac{ 1 }{ n ^{6} }1\] It just didn't seem right to me when I wrote it down.

OpenStudy (asnaseer):

it is correct but you can simplify this further (e.g. anything times 1 is just the same result)

OpenStudy (asnaseer):

\[m^3n^{-6}p^0=m^3\frac{1}{n^6}1=\frac{m^3}{n^6}\]

OpenStudy (asnaseer):

does that make sense to you?

OpenStudy (anonymous):

It does now. Thanks for helping. Trying to get my high school diploma and math was my worst subject in school.

OpenStudy (asnaseer):

yw :) and I am certain you will pick this up quickly - you certainly have the right attitude towards it :)

OpenStudy (anonymous):

Thanks. Hope can remain so or else I will have no hair left. It really is the letters the throw me off the most though

OpenStudy (asnaseer):

ha ha - I see you have a good sense of humour as well :) good luck with your diploma my friend! :)

OpenStudy (asnaseer):

This is also a very good site with lots of short videos and exercises that may help you: https://www.khanacademy.org/math/algebra/exponent-equations/exponent-properties-algebra/e/simplifying_expressions_with_exponents

OpenStudy (anonymous):

Oh thank you. Every little bit helps. I would give a medal but have no idea how.

OpenStudy (asnaseer):

I am here to help rather than to collect medals :) However, the way you medal someone is by clicking on the "Best Response" button next to ay of their replies.

OpenStudy (asnaseer):

BTW: Once you are happy that your question has been answered then you should Close it.

OpenStudy (anonymous):

Oh, Sorry, Will close now. Thanks for the help. Happy Holidays.

OpenStudy (asnaseer):

thx :)